# Negative non-Gaussianity as a salvager for PBHs with PTAs in bounce

### v1

@@ II Bounce from EFT framework
Our next objective is to construct the scalar power spectrum based on the mode solutions of the MS Eqn.(9) across all the regions of interest within the framework. These regions include the contraction phase followed by a bouncing phase that transitions to the inflationary stages of the first slow-roll (SRI), an ultra slow-roll (USR), and the final phase of inflation with a second slow-roll (SRII). The detailed structure of the curvature perturbation mode solutions is elaborated in the appendix A. For an in-depth derivation, we direct the reader to [240, 241, 303]. Here, we focus on the structure of the scalar power spectrum and subsequent contributions from quantum loop effects that yield us the final power spectrum.
@@ III.1 Tree Level Contribution
We need to employ the Fourier space solutions for the curvature perturbation in each phase to construct the full tree-level power spectrum as per the above definition. The resulting version from each phase is combined to give us the late-time version $(\tau_{0}\rightarrow 0)$ as follows [240, 241, 303]:
@@
where the top line introduces contributions from the SRI phase, contraction, and bouncing phase, respectively, while the bottom line shows the contributions from the USR and the SRII phases. The quantities $\{\alpha_{2},\beta_{2},\alpha_{3},\beta_{3}\}$ represent the Bogoliubov coefficients that appears as we pass through each phase in the inflationary stages. These coefficients signal a change in the underlying vacuum structure from the initial conditions at the start of inflation. Notice that the transition from contraction to the bounce and into the SRI is continuous such that the underlying choice of quantum vacuum structure remains the same as the initial Bunch-Davies vacuum. However, we require the implementation of a Heaviside theta function, $\Theta(k-k_{i}),\;i=\{s,e\}$, which indicates the sharp transition criteria leading to a shift from the initial vacuum structure. Consequently, it is only when we begin to see transitions of the SRI-USR $(k=k_{s})$ and USR-SRII $(k=k_{e})$, that we need to incorporate the modified quantum vacuum. The wavenumber $k_{*}$ refer to the reference scale which we set here to be the pivot scale, $k_{*}=0.02\;{\rm Mpc^{-1}}$. Also important are the quantities, $\epsilon_{c},\epsilon_{b}$, that refer to the choices of the first SR parameter in the contraction and bouncing phases, respectively, and the choice signifies the particular nature whether it is an ekpyrotic kind or matter kind of phase.
@@ III.2 Quantum Loop Corrected
We now examine the quantum loop corrections to the tree-level power spectrum structure from each phase within our framework discussed in the previous section. The task here requires expanding the EFT action to obtain the cubic order curvature perturbation action and working with the Schwinger-Keldysh or in-in formalism to evaluate the one-loop corrections to the previously mentioned $2$-point function. In the appendix B, we provide the necessary details concerning the computations and this formalism to the reader. Here, we provide only the relevant expressions that, when combined, give us the final quantum loop corrected version of the scalar power spectrum. To present the procedure more succinctly, we divide it into two steps and explain the consequences of each outcome.
@@ III.2.1 $\diamond$ Regularization
The primary task with all the relevant one-loop correlation functions, using the terms from the cubic order action, would be to handle the ultraviolet (UV) and infrared (IR) divergences. The initial technique employed for this purpose is the cut-off regularization, which is to compute the necessary loop integrals. These consists of a finite wavenumber interval for the momentum integration, preferably the modes of significance of the phase of interest, while for the temporal integration the late time limit $(\tau_{0}\rightarrow 0)$ is opted later to gather information at the super-horizon scales. Although we do not perform the said computations explicitly, we do provide the final expressions resulting from these in the appendix C. After evaluating the loop integrals, the cut-off regularized one-loop scalar power spectrum can be written as follows:
@@
The quantity in Eqn. (16) becomes our first non-trivial result after controlling the presence of divergences from quantum loops. An immediate consequence of such a regularization procedure, however, is the presence of regularization scheme-dependent counter-terms that must be adjusted to remove UV divergent contributions, and this is the next topic of discussion.
@@ III.2.2 $\diamond$ Renormalization
where the now regularized-renormalized one-loop integrals give us the terms, $\overline{{\bf T}}_{\bf C}$ (contraction), $\overline{{\bf T}}_{\bf B}$ (bounce), $\overline{{\bf T}}_{\bf SRI}$ (SRI), $\overline{{\bf T}}_{\bf USR}$ (USR), and $\overline{{\bf T}}_{\bf SRII}$ (SRII), and their explicit forms are given in appendix C.
The leftover IR divergences yet remain irremovable and require further analysis. The first step in this direction is the power spectrum renormalization. This technique helps us control/smooth the IR-divergent corrections of the power spectrum induced from all the pre-inflationary (contraction and bouncing) phases and the SRI, USR, and SRII phases. To execute this technique, we introduce an appropriate IR counter-term to renormalize the correlation functions. The estimation of this counter-term depends on our UV counter-terms, evaluated previously to renormalize the power-law divergences, and a renormalization condition accepted at the CMB pivot scale $(k_{*})$. For proper details regarding the various steps in the calculations, we refer the reader to [240]. We now show the consequence of the above explained steps which modify our previous result via the IR counter-term as follows:
@@
We notice the use of similar terms, as before in Eqn. (19), but this time with them getting evaluated at the pivot scale, hence the asterisk, as per the CMB renormalization condition. After performing the product between the various terms, we get the important result where the linear order term is completely devoid of any logarithmic IR corrections, and the remaining divergences are pushed to higher loop orders in the perturbative expansion. In the next section, we discuss the strategy to resum such contributions.
@@ III.2.3 $\diamond$ Dynamical RG Resummation
We begin from the results of previous discussions on the power spectrum renormalization technique that was able to soften the IR divergences by pushing them to the higher loop orders and removing their presence at the linear order. Eventually, we must take care of all the large logarithms, which display a disturbing feature of secular growth with time, to extract the well-needed behavior of these correlations at late times. To complete this task of producing a finite result by handling the divergent contributions carefully, we soften the remaining logarithmic IR divergences and carry out their resummation with the help of the dynamical renormalization group (DRG) technique [441, 442, 443, 444, 445, 446, 447, 448, 310, 449, 450]. The primary advantage of DRG here is the ability to resum the softened logarithmic terms by absorbing the divergent or secularly growing terms coming at all orders in the perturbative expansion. At the end we get a finite, convergent result, that is resummed at all loop orders. We refer the reader to the [240] for the full details on calculations from this procedure and to the appendix C for the explicit list of expressions used here. Here, we mention the final expression that remains useful for the latter sections of this paper, where we deal with PBH formation, [240]:
@@
and which involves the pivot scale amplitude defined before in Eqn. (18). We note some crucial facts about the DRG analysis here before moving with the latter part of this paper. The foremost outcome of the DRG resummed version of the one-loop corrected power spectrum is that, unlike in the renormalized one-loop power spectrum, the two-point function is reliable without any problems at large times from the uncontrolled growth of amplitude due to IR divergent logarithms. The final amplitude gets exponentiated which is a crucial outcome from this procedure in general and this leads to its rapid fall off instead of the growth. Another fact is that, in the computations, we do not demand any exact details of the higher-order loop diagrams during resummation neither it needs the dominance of all chain diagrams over other possible diagrams in the computation, but it will definitely increase the leading contributions from these components with Logarithms.
@@ IV Primordial Black Hole Formation
In the previous section, we presented the development of the regularized-renormalized-resummed scalar power spectrum where we carefully handled the quantum one-loop divergences that lead to a finite outcome. This result can now be utilized in the subsequent sections to investigate the generation of PBHs. The standard theory behind PBH formation begins with the journey of curvature perturbations stretched on the super-horizon scales where they freeze in amplitude and, as inflation ends, these fluctuations start to re-enter the Hubble horizon. Upon their re-entry, these fluctuations leave their signatures in the primordial plasma that appears in the form of density fluctuations, and, with time, they develop regions of densities larger or smaller than the mean background density of the universe. Due to gravitational attraction, if these overdense regions cross a certain threshold value, they suffer gravitational collapse, leading to the generation of PBH. The standard theory was built considering such density fluctuations in the radiation-dominated (RD) era where we have the relation, $c_{s}^{2}=w=1/3$, between the sound speed and equation of state of this era. The exact mechanism of the collapse of perturbations into PBHs has been a topic of intense study for the past few decades. There have been numerous mechanisms exploring this problem, the most widely investigated of those include the Press-Schechter theory [290, 125], the peak theory [451, 452, 453, 454, 455, 456], and the compaction function approach [457, 458, 459, 460, 461, 462, 463, 464, 305, 465, 306, 28, 307, 23, 280, 287, 466]. In the following sections, we will focus on the details of the compaction function approach and use it as a foundation for our analysis to address the problems related to PBHs production.
@@ IV.2.2 $\diamond$ Distribution and Threshold Statistics with Compaction Function
where we observe the cumulative contribution to the correlations coming after integration over each phase. Notably, the USR phase $(k_{s},k_{e})$ yields the largest contribution attributable to the significant amplification of the scalar power spectrum during this interval. Contrary to the scales corresponding to SRI or SRII, this is the scale where the curvature peturbation attains substantial amplification for generating the PBHs.
$\int_{0}^{\infty}\equiv\underbrace{\int_{{k}_{c}}^{k_{s}}}_{\textbf{negligible}}+\underbrace{\int_{{k}_{s}}^{k_{e}}}_{\textbf{Dominant(USR)}}+\underbrace{\int_{{k}_{e}}^{k_{end}}}_{\textbf{negligible}}$ (61)
The cumulative impact of the chosen window and transfer functions through these correlators can be visualized in the PDF later in the numerical analysis section. We now take a moment to comment on the behaviour of the joint PDF on the basis of the possible correlations.
@@
where $M_{\odot}$ is the solar mass, $\Omega_{\rm DM}=0.674$ is the current dark matter density. We will analyze this quantity in the coming sections, discussing various related numerical outcomes from our study by incorporating the PNGs.
@@ VI.3 PBH Overproduction
With this consideration in mind, we return to the analysis of figure (2). Notably, all ${\cal C}_{\rm th}$ values share one common boundary denoted by the dotted black line. The iso-contours lines for $f_{\rm NL}=-39.95$ (figure 2) show a pronounced overlap of the domain with the PDF for lower masses, indicating a greater probability of obtaining sizeable abundance for these lower mass PBHs than for the higher mass. A parallel trend is evident for $f_{\rm NL}=-35/8$ (figure 2), where the domain gets drawn inwards for the low masses. However, a direct comparison between the two $f_{\rm NL}$ values reveals that contour lines for $f_{\rm NL}=-39.95$ exhibit a greater compatibility with the domain within the distribution. This implies that $f_{\rm NL}=-39.95$ is more favourable for the scenario of ample PBHs production as compared to $f_{\rm NL}=-35/8$. The above observation is further corroborated by the illustrations in the plots of figure (3), which demonstrates the influence of NGs on the power spectrum, varying along with the transition wavenumber $k_{s}$ in the USR region, necessary to obtain a suitable range of fractional abundance and subsequently avoid overproduction. The results are correlated with the PTA data to seek observational constraints. The various coloured bands represent the PBHs abundance in the range $f_{\rm PBH}\in(10^{-3}-1)$ corresponding to thresholds ${\cal C}_{\rm th}\in(0.5,0.6,0.65)$. Before diving into the discussion of the figure, we like to present a detail about the suitability of our forthcoming analysis with the NANOGrav15 and EPTA posteriors. A careful examination will display a resemblance of our chosen scalar power spectrum with the log-normal power spectrum. It is indebted to the presence of a USR phase with a sharp peak. This resemblance justifies our use of the PTA log-normal based posteriors for analysis. Additionally, the SIGW spectrum displayed in fig.(1) also exhibits a causality tail in the IR region, which is a distinctive feature when using the log-normal power spectrum. This further reinforces the idea that a separate probability analysis is not required in our case to develop the posteriors for deriving our outcomes.
@@
Now, in figure 3 with $f_{\rm NL}=-39.95$, and $c_{s}=1$, we observe that all the bands corresponding to the threshold values yield substantial abundance with amplitudes in the range of $A\in(10^{-1}-10^{-0.5})$. Consequently, they effectively circumvent the overproduction problem by residing within the $1\sigma$ region of the NANOGrav15 posteriors. Conversely, reducing the magnitude of the negative $f_{\rm NL}$ to $-35/8$, and keeping $c_{s}=1$, would alter the results as shown in figure 3. Here for the threshold ${\cal C}_{\rm th}=0.65$ (brown band), both the upper and lower limits of the $f_{\rm PBH}$ fall within the $1\sigma$ region of the NANOGrav15 and EPTA contours, while for ${\cal C}_{\rm th}=0.6$ (orange band), the upper value of $f_{\rm PBH}$ merely grazes from outside of the $1\sigma$ contour of the EPTA. Additionally, for ${\cal C}_{\rm th}=0.5$ (purple band) in figure 3, the result is further constrained by being excluded at over the $1\sigma$ contour of the EPTA and barely grazes the $1\sigma$ region of the NANOGrav15. This observation aligns well with the theoretical expectations since the perturbation amplitude is intricately linked with the compaction threshold, and a higher ${\cal C}_{\rm th}$ within the linear regime necessitates a greater amplitude to generate the same fractional abundance as compared to lower values of the same. A similar trend is visible if we take $c_{s}=0.88$ and examine their results from the figs. (3, 3) for the two distinct $f_{\rm NL}$ cases. However, if we instead draw a direct comparison to inspect the impact of the effective sound speed $c_{s}$ on the amplitude, we can readily witness a significant difference separately captured in figure 4.
@@
The left 4 and right 4 figures represent the variation of the power spectra amplitude with $k_{s}$ for $c_{s}=0.88$, and $c_{s}=1$, respectively. The plots have been realized with ${\cal C}_{\rm th}=0.65$. Figure 4 demonstrates both the bands, i.e. the orange for $f_{\rm NL}=-39.95$, and purple for $f_{\rm NL}=-35/8$ remain largely excluded over the $1\sigma$ region of the PTA contours. While the orange band circumvents overproduction and marginally favours the $2\sigma$ region of the NANOGrav15 contour, the scenario with $f_{\rm NL}=-35/8$ fails to generate sufficient amplitude to achieve comparable abundance while simultaneously remaining consistent with the NANOGrav15. Attempts to amplify the amplitude to conform to the SIGW interpretation of the NANOGrav15 data would bring the risk of overproducing PBHs. Now, in contrast, figure 4 temptingly highlights a dramatic surge in the amplitude for both the $f_{\rm NL}$ values. This remarkable shift prompted by the alteration of $c_{s}$ underscores the sensitivity of the setup to the effective sound speed parameter. Lowering $c_{s}$ from $1$ to $0.88$ drops the power spectrum amplitude to $A=10^{-3}$, an order of magnitude less when compared to $c_{s}=1$. This reduction in amplitude is consequential because it directly impacts the conditions requisite for PBH formation, as amplitude is integrated into the $f_{\rm PBH}$ through the Eqn. (72). The likelihood of PBH production drops further, drastically dropping for choices below $c_{s}=0.88$, making such scenarios irrelevant for analysis. Moreover, restraining to $c_{s}\leqslant 1$ is essential to uphold the causality constraints. Exceeding $c_{s}=1$ would jeopardize the perturbativity conditions and therefore can be ruled out.
@@
Additionally, we notice the appearance of features in the width of the bands of figure (3). Notably, they are not uniform across the transition wavenumbers $k_{s}$. These bands are seen to exhibit greater thickness near $k_{s}\sim{\cal O}(10^{6}{\rm Mpc^{-1}})$ and become narrower towards the right side of the plots. Note that the transition wavenumber is linked with the PBHs mass by Eqn. (63). This is indicative of the idea that with increased wavenumber, PBHs with lower masses are poised to produce substantial abundance with a slightly lesser amplitude as compared to those with higher masses.
@@
We now examine the results for the fractional abundance of PBHs represented by the figure (5). The three inclined coloured lines on each of these plots correspond to distinct compaction thresholds ${\cal C}_{\rm th}=0.65$ (orange), ${\cal C}_{\rm th}=0.6$ (purple), and ${\cal C}_{\rm th}=0.5$ (green). These lines depict the variation of the PBH abundance with the regularized-renormalized-resummed power spectrum amplitude for the two principal values of $f_{\rm NL}$ and $c_{s}$, recurrently mentioned in this paper. The transition wavenumber is kept fixed here at $k_{s}=10^{7}{\rm Mpc^{-1}}$. As clearly evident from the graphs, opting for a higher value of ${\cal C}_{\rm th}$ calls for a larger amplitude across all the generated plots, and vice versa. Further, to obtain a suitable $f_{\rm PBH}$ within the range $(10^{-3},1)$ with $(f_{\rm NL},c_{s})=(-39.95,1)$, we require a peak amplitude $A\sim{\cal O}(0.1)$. The abundance is extremely susceptible to the change in peak amplitude irrespective of the chosen threshold condition and any specific combination of the two $f_{\rm NL}$ and $c_{s}$ values. The scenario with $f_{\rm NL}=-35/8$ and $c_{s}=0.88$ (non-canonical) leads to the minimal amount of peak amplitude, while the negatively large $f_{\rm NL}=-39.95$ and $c_{s}=1$ (canonical) scenario leads to a huge amplification of $A$ which comes extremely close to breaching the perturbativity conditions. Now, having presented the results of our setup, let us look at how it compares to alternative frameworks and models previously studied to address the recent crucial overproduction problem.

### v3

@@ II Bounce from EFT framework
Our next objective is to construct the scalar power spectrum based on the mode solutions of the MS Eqn.(9) across all the regions of interest within the framework. These regions include the contraction phase followed by a bouncing phase that transitions to the inflationary stages of the first slow-roll (SRI), an ultra slow-roll (USR), and the final phase of inflation with a second slow-roll (SRII). The detailed structure of the solutions of the curvature perturbation mode is elaborated in the Appendix B. For an in-depth derivation, we direct the reader to [253, 254, 319]. Here, we focus on the structure of the scalar power spectrum and subsequent contributions from quantum loop effects that yield the final power spectrum.
@@ III.1 Tree Level Contribution
We need to employ the Fourier space solutions for the curvature perturbation in each phase to construct the full tree-level power spectrum as per the above definition. The resulting version coming from each phase is combined to give us the late-time version $(\tau_{0}\rightarrow 0)$ as follows [253, 254, 319]:
@@
where the top line introduces contributions from the SRI phase, contraction, and bouncing phase, respectively, while the bottom line shows the contributions from the USR and SRII phases. The quantities $\{\alpha_{2},\beta_{2},\alpha_{3},\beta_{3}\}$ represent the Bogoliubov coefficients that appear as we pass through each phase in the inflationary stages. These coefficients signal a change in the underlying vacuum structure from the initial conditions at the start of inflation. Notice that the transition from contraction to the bounce and into the SRI is continuous such that the underlying choice of quantum vacuum structure remains the same as the initial Bunch-Davies vacuum. However, we require the implementation of a Heaviside theta function, $\Theta(k-k_{i}),\;i=\{s,e\}$, which indicates the sharp transition criteria leading to a shift from the initial vacuum structure. Consequently, it is only when we begin to see transitions of the SRI-USR $(k=k_{s})$ and USR-SRII $(k=k_{e})$, that we need to incorporate the modified quantum vacuum. The wavenumber $k_{*}$ refers to the reference scale which we set here as the pivot scale, $k_{*}=0.02\;{\rm Mpc^{-1}}$. Also important are the quantities, $\epsilon_{c},\epsilon_{b}$, which refer to the choices of the first SR parameter in the contraction and bouncing phases, respectively, and the choice signifies the particular nature whether it is an ekpyrotic kind or a matter kind of phase.
@@ III.2 Quantum Loop Corrected
We now examine the quantum loop corrections to the tree-level power spectrum structure from each phase within our framework discussed in the previous section. The task here requires expanding the EFT action to obtain the cubic-order curvature perturbation action and working with the Schwinger-Keldysh or in-in formalism to evaluate the one-loop corrections to the previously mentioned $2$-point function. In the Appendix C, we provide the necessary details concerning the computations and this formalism to the reader. Here, we provide only the relevant expressions that, when combined, give us the final quantum loop corrected version of the scalar power spectrum. To present the procedure in a succinct manner, we divide it into two steps and explain the consequences of each outcome.
@@ III.2.1 $\diamond$ Regularization
The primary task with all the relevant one-loop correlation functions, using the terms from the cubic-order action, would be to handle the ultraviolet (UV) and infrared (IR) divergences. The initial technique employed for this purpose is the cut-off regularization, which is to compute the necessary loop integrals. These consist of a finite wavenumber interval for the momentum integration, preferably the modes of significance of the phase of interest, while for the temporal integration the late time limit $(\tau_{0}\rightarrow 0)$ is opted later to gather information at the super-horizon scales. Although we do not perform the said computations explicitly, we do provide the final expressions resulting from these in the Appendix D. After evaluating the loop integrals, the cut-off regularized one-loop scalar power spectrum can be written as follows:
@@
The quantity in Eqn. (16) becomes our first non-trivial result after controlling the presence of divergences from quantum loops. An immediate consequence of such a regularization procedure, however, is the presence of regularization scheme-dependent counterterms that must be adjusted to remove UV divergent contributions, and this is the next topic of discussion.
@@ III.2.2 $\diamond$ Renormalization
where the now regularized-renormalized one-loop integrals give us the terms, $\overline{{\bf T}}_{\bf C}$ (contraction), $\overline{{\bf T}}_{\bf B}$ (bounce), $\overline{{\bf T}}_{\bf SRI}$ (SRI), $\overline{{\bf T}}_{\bf USR}$ (USR), and $\overline{{\bf T}}_{\bf SRII}$ (SRII), and their explicit forms are given in Appendix D.
The leftover IR divergences remain irremovable and require further analysis. The first step in this direction is the power spectrum renormalization. This technique helps us control/smooth the IR-divergent corrections of the power spectrum induced from all the pre-inflationary (contraction and bouncing) phases and the SRI, USR, and SRII phases. To execute this technique, we introduce an appropriate IR counter-term to renormalize the correlation functions. The estimation of this counterterm depends on our UV counterterms, evaluated previously to renormalize the power-law divergences, and a renormalization condition accepted at the CMB pivot scale $(k_{*})$. For proper details regarding the various steps in the calculations, we refer the reader to [253]. We now show the consequence of the above explained steps which modify our previous result via the IR counter-term as follows:
@@
We notice the use of similar terms, as before in Eqn. (19), but this time with them evaluated at the pivot scale, hence the asterisk, as per the CMB renormalization condition. After performing the product between the various terms, we get the important result where the linear-order term is completely devoid of any logarithmic IR corrections, and the remaining divergences are pushed to higher loop orders in the perturbative expansion. In the next section, we discuss the strategy for resumming such contributions.
@@ III.2.3 $\diamond$ Dynamical RG Resummation
We begin from the results of previous discussions on the power spectrum renormalization technique, which was able to soften the IR divergences by pushing them to the higher loop orders and removing their presence at the linear order. Eventually, we must take care of all the large logarithms, which display a disturbing feature of secular growth with time, to extract the well-needed behavior of these correlations at late times. To complete this task of producing a finite result by carefully handling divergent contributions, we soften the remaining logarithmic IR divergences and perfrom their resummation with the help of the dynamical renormalization group (DRG) technique [457, 458, 459, 460, 461, 462, 463, 464, 326, 465, 466]. The primary advantage of DRG here is the ability to resum the softened logarithmic terms by absorbing the divergent or secularly growing terms coming at all orders in the perturbative expansion. At the end we get a finite, convergent result that is resummed at all loop orders. We refer the reader to [253] for the full details on calculations from this procedure and to the Appendix D for the explicit list of expressions used here. Here, we mention the final expression that remains useful for the latter sections of this paper, where we deal with PBH formation, [253]:
@@
and which involves the pivot scale amplitude defined before in Eqn. (18). We note some crucial facts about the DRG analysis here before moving on to the latter part of this paper. The foremost outcome of the DRG resummed version of the one-loop corrected power spectrum is that, unlike in the renormalized one-loop power spectrum, the two-point function is reliable without any problems at large times from the uncontrolled growth of amplitude due to IR divergent logarithms. The final amplitude gets exponentiated, which is a crucial outcome of this procedure in general, and this leads to its rapid fall off instead of the growth. Another fact is that, in the computations, we do not demand any exact details of the higher-order loop diagrams during resummation nor it needs the dominance of all chain diagrams over other possible diagrams in the computation, but it will definitely increase the leading contributions from these components with Logarithms.
@@ IV Primordial Black Hole Formation
In the previous section, we presented the development of the regularized-renormalized-resummed scalar power spectrum, where we carefully handled the quantum one-loop divergences that lead to a finite outcome. This result can now be utilized in the subsequent sections to investigate the generation of PBHs. The standard theory behind PBH formation begins with the journey of curvature perturbations stretched on the super-horizon scales where they freeze in amplitude and, as inflation ends, these fluctuations start to re-enter the Hubble horizon. Upon their re-entry, these fluctuations leave their signatures in the primordial plasma that appears in the form of density fluctuations, and, with time, they develop regions of densities larger or smaller than the mean background density of the universe. Due to gravitational attraction, if these overdense regions cross a certain threshold value, they suffer gravitational collapse, leading to the generation of PBH. The standard theory was built considering such density fluctuations in the radiation-dominated (RD) era, where we have the relation, $c_{s}^{2}=w=1/3$, between the sound speed and the equation of state of this era. The exact mechanism of the collapse of perturbations into PBHs has been a topic of intense study for the past few decades. There have been numerous mechanisms exploring this problem, the most widely investigated of them include the Press-Schechter theory [306, 136], the peak theory [467, 468, 469, 470, 471, 472], and the compaction function approach [473, 474, 475, 476, 477, 478, 479, 480, 321, 481, 322, 28, 323, 23, 296, 303, 482]. In the following sections, we will focus on the details of the compaction function approach and use it as a foundation for our analysis to address the problems related to PBHs production.
@@ IV.2.2 $\diamond$ Distribution and Threshold Statistics with Compaction Function
where we observe the cumulative contribution to the correlations coming after integration over each phase. Notably, the USR phase $(k_{s},k_{e})$ yields the largest contribution attributable to the significant amplification of the scalar power spectrum during this interval. Contrary to the modes corresponding to either contraction $(k_{c}<k<k_{b})$, bounce $(k_{b}<k<k_{*})$, SRI $(k_{*}<k<k_{s})$, or SRII $(k_{e}<k<k_{\rm end})$, it is for those scales inside the USR where the curvature perturbation attains substantial amplification for generating the PBHs,
$\int_{0}^{\infty}\equiv\underbrace{\int_{0}^{k_{c}}}_{\textbf{negligible}}+\underbrace{\int_{{k}_{c}}^{k_{b}}}_{\textbf{negligible}}+\underbrace{\int_{{k}_{b}}^{k_{*}}}_{\textbf{negligible}}+\underbrace{\int_{{k}_{*}}^{k_{s}}}_{\textbf{suppressed}}+\underbrace{\int_{{k}_{s}}^{k_{e}}}_{\textbf{Dominant(USR)}}+\underbrace{\int_{{k}_{e}}^{k_{\rm end}}}_{\textbf{suppressed}}+\underbrace{\int_{k_{\rm end}}^{\infty}}_{\textbf{suppressed}}$ (61)
where the contributions keep getting more negligible as we consider modes from before the inflationary phases. The cumulative impact of the chosen window and transfer functions through these correlators can be visualized in the PDF later in the numerical analysis section. We now take a moment to comment on the behaviour of the joint PDF on the basis of the possible correlations.
@@
where $M_{\odot}$ is the solar mass, $\Omega_{\rm DM}=0.264$ is the current dark matter density. We will analyze this quantity in the coming sections, discussing various related numerical outcomes from our study by incorporating the PNGs.
@@ VI.3 PBH Overproduction
With this consideration in mind, we return to the analysis of figure (2). Notably, all ${\cal C}_{\rm th}$ values share one common boundary, denoted by the dotted black line. The iso-contour lines for $f_{\rm NL}=-39.95$ (figure 2) show a pronounced overlap of the domain with the PDF for lower masses, indicating a greater probability of obtaining a sizeable abundance for these lower mass PBHs than for the higher mass. A parallel trend is evident for $f_{\rm NL}=-35/8$ (figure 2), where the domain gets drawn inwards for the low masses. However, a direct comparison between the two $f_{\rm NL}$ values reveals that the contour lines for $f_{\rm NL}=-39.95$ exhibit a greater compatibility with the domain within the distribution. This implies that $f_{\rm NL}=-39.95$ is more favourable for the scenario of ample PBHs production as compared to $f_{\rm NL}=-35/8$. The above observation is further corroborated by the illustrations in the plots of figure (3), which demonstrates the influence of NGs on the power spectrum, which varies along with the transition wavenumber $k_{s}$ in the USR region, necessary to obtain a suitable fractional abundance range and subsequently avoid overproduction. The results are correlated with the PTA data to find observational constraints. The various coloured bands represent the PBHs abundance in the range $f_{\rm PBH}\in(10^{-3}-1)$ corresponding to thresholds ${\cal C}_{\rm th}\in(0.5,0.6,0.65)$. Before diving into the discussion of the figure, we like to present a detail about the suitability of our forthcoming analysis with the NANOGrav15 and EPTA posteriors. A careful examination will display a resemblance of our chosen scalar power spectrum with the log-normal power spectrum. This is indebted to the presence of a USR phase with a sharp peak. This resemblance justifies our use of the PTA log-normal-based posteriors for analysis. Additionally, the SIGW spectrum displayed in fig.(1) also exhibits a causality tail in the IR region, which is a distinctive feature when using the log-normal power spectrum. This further reinforces the idea that separate probability analysis is not required in our case to develop posteriors for deriving our outcomes.
@@
Now, in figure 3 with $f_{\rm NL}=-39.95$, and $c_{s}=1$, we observe that all the bands corresponding to the threshold values yield substantial abundance with amplitudes in the range of $A\in(10^{-1}-10^{-0.5})$. Consequently, they effectively circumvent the overproduction problem by residing within the $1\sigma$ region of the NANOGrav15 posteriors. Conversely, reducing the magnitude of the negative $f_{\rm NL}$ to $-35/8$, and keeping $c_{s}=1$, would alter the results as shown in figure 3. Here for the threshold ${\cal C}_{\rm th}=0.65$ (brown band), both the upper and lower limits of $f_{\rm PBH}$ fall within the $1\sigma$ region of the NANOGrav15 and EPTA contours, while for ${\cal C}_{\rm th}=0.6$ (orange band), the upper value of $f_{\rm PBH}$ merely grazes from outside of the $1\sigma$ contour of the EPTA. Additionally, for ${\cal C}_{\rm th}=0.5$ (purple band) in figure 3, the result is further constrained by being excluded over the $1\sigma$ contour of the EPTA and barely grazes the $1\sigma$ region of NANOGrav15. This observation aligns well with theoretical expectations since the perturbation amplitude is intricately linked with the compaction threshold, and a higher ${\cal C}_{\rm th}$ within the linear regime necessitates a greater amplitude to generate the same fractional abundance as compared to lower values of the same. A similar trend is visible if we take $c_{s}=0.88$ and examine their results from the figs. (3, 3) for the two distinct $f_{\rm NL}$ cases. However, if we instead draw a direct comparison to inspect the impact of the effective sound speed $c_{s}$ on the amplitude, we can easily witness a significant difference separately captured in figure 4.
@@
The left 4 and right 4 figures represent the variation of the power spectra amplitude with $k_{s}$ for $c_{s}=0.88$, and $c_{s}=1$, respectively. The plots have been realized with ${\cal C}_{\rm th}=0.65$. Figure 4 shows both the bands, i.e. the orange for $f_{\rm NL}=-39.95$, and purple for $f_{\rm NL}=-35/8$ remain largely excluded over the $1\sigma$ region of the PTA contours. Although the orange band circumvents overproduction and marginally favours the $2\sigma$ region of the NANOGrav15 contour, the scenario with $f_{\rm NL}=-35/8$ fails to generate sufficient amplitude to achieve a comparable abundance while simultaneously remaining consistent with the NANOGrav15. Attempts to amplify the amplitude to conform to the SIGW interpretation of the NANOGrav15 data would bring the risk of overproducing PBHs. Now, in contrast, figure 4 temptingly highlights a dramatic surge in the amplitude for both the $f_{\rm NL}$ values. This remarkable shift prompted by the alteration of $c_{s}$ underscores the sensitivity of the setup to the effective sound speed parameter. Lowering $c_{s}$ from $1$ to $0.88$ drops the power spectrum amplitude to $A=10^{-3}$, an order of magnitude less when compared to $c_{s}=1$. This reduction in amplitude is consequential because it directly impacts the conditions requisite for PBH formation, as amplitude is integrated into the $f_{\rm PBH}$ through the Eqn. (72). The likelihood of PBH production drops further, drastically dropping for choices below $c_{s}=0.88$, making such scenarios irrelevant for analysis. Moreover, restraining to $c_{s}\leqslant 1$ is essential to uphold causality constraints. Exceeding $c_{s}=1$ would jeopardize the perturbativity conditions and therefore can be ruled out.
@@
Additionally, we notice the appearance of features in the width of the bands in the figure (3). Notably, they are not uniform across the transition wavenumbers $k_{s}$. These bands are seen to exhibit a greater thickness near $k_{s}\sim{\cal O}(10^{6}{\rm Mpc^{-1}})$ and become narrower toward the right side of the plots. Note that the transition wavenumber is linked with the PBHs mass by Eqn. (63). This is indicative of the idea that with increased wavenumber, PBHs with lower masses are poised to produce substantial abundance with a slightly lower amplitude as compared to those with higher masses.
@@
We now examine the results for the fractional abundance of PBHs represented by the figure (5). The three inclined coloured lines on each of these plots correspond to distinct compaction thresholds ${\cal C}_{\rm th}=0.65$ (orange), ${\cal C}_{\rm th}=0.6$ (purple), and ${\cal C}_{\rm th}=0.5$ (green). These lines show the variation of the PBH abundance with the regularized-renormalized-resummed power spectrum amplitude for the two principal values of $f_{\rm NL}$ and $c_{s}$, recurrently mentioned in this paper. The transition wavenumber is kept fixed here at $k_{s}=10^{7}{\rm Mpc^{-1}}$. As clearly evident from the graphs, opting for a higher value of ${\cal C}_{\rm th}$ calls for a larger amplitude across all the generated plots, and vice versa. Further, to obtain a suitable $f_{\rm PBH}$ within the range $(10^{-3},1)$ with $(f_{\rm NL},c_{s})=(-39.95,1)$, we require a peak amplitude $A\sim{\cal O}(0.1)$. The abundance is extremely susceptible to the change in peak amplitude irrespective of the chosen threshold condition and any specific combination of the two $f_{\rm NL}$ and $c_{s}$ values. The scenario with $f_{\rm NL}=-35/8$ and $c_{s}=0.88$ (non-canonical) leads to the minimal amount of peak amplitude, while the negatively large $f_{\rm NL}=-39.95$ and $c_{s}=1$ (canonical) scenario leads to a huge amplification of $A$ that comes extremely close to breaching the perturbativity conditions. Now, having presented the results of our setup, let us look at how it compares to alternative frameworks and models previously studied to address the recent crucial overproduction problem.